Given the quadratic function f(x) = 9x^2 + 36x + 36, find a value of x such that f(x) = 9. Can you please show the work too?
Thanks in advance!

Found 2 solutions by ewatrrr, solver91311:
Answer by ewatrrr(24785)


Hi
f(x) = 9x^2 + 36x + 36
f(x) = 9
9 = 9x^2 + 36x + 36
9x^2 + 36c + 27 = 0
3x^2 + 12c + 9 = 0
factoring
(3x + 3)(x + 3)= 0 Note:SUM of the inner product(3x) and the outer product(9x) = 12x
(3x + 3)=0
x = -1
(x + 3)= 0
x = -3

Answer by solver91311(24713)



If



AND



Then



Which is to say:



Just solve the quadratic using the quadratic formula.


John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism